# GenKI: Enhancing Open-Domain Question Answering with Knowledge Integration and Controllable Generation in Large Language Models

### v1

@@ 4. Methodology
To address the knowledge deficiency and uncontrollable generation issues in LLMs for OpenQA, our system combines three modules as illustrated in Figure 2. On the whole, we decompose the model $A=M(q,F,P)$ into $A=C(A_{K},F)=C(K(q,P),F)$ is to ensure that the model focuses on either knowledge integration or controlled generation in each instance, to avoid the distributional differences brought by these two tasks. Specifically, we first train a Knowledge Retriever module $R(P,q)=P_{R}$ to retrieve knowledge according to the query supplying the model during knowledge integration steps. Subsequently, once the retrieved knowledge is synthesized, a Knowledge Integration module $K(P_{R},q)=A_{K}$ (Part A in Figure  2) is deployed to generate answers by integrating the retrieved knowledge. Finally, a Controllable Generation module $C(A_{K},F)=A$ (Part B in Figure 2) involves using a fine-tuned LLM on the target dataset format for post-processing the results and a Reward model to further ensure that the output of our method is controlled. We split $C(A_{K},F)=Ensemble(Rm(A_{K},F),Ext(A_{K},F))$ (Part C in Figure 2) to gain formatted, coherent and factual answer. The function of each module will be introduced in detail in the following subsections.
Figure 2. The overall framework of GenKI. The left part presents several examples of our prompts used in the knowledge integration and controllable generation module. The right part outlines the process of tuning LLMs to learn knowledge and answer questions in specific task situations.
@@ 4.1. Knowledge Retriever
Firstly, given vast knowledge repositories, we should prioritize enabling our model to learn from them. However, feeding all the sentences $P$ from the knowledge base into the LLM may potentially result in the model’s inability to grasp essential focal points. Therefore, we need a retriever $R(P,q)=P_{R}\subset P$ as the initial input provided for the next knowledge integration module $K$.
Specifically, we use a DPR to retrieve knowledge from the knowledge base $P$ according to the question. Setting the encoder as $f$, and the trainable parameters as $\Phi$, DPR computes question-sentence similarity for a given question q and a set of sentences P using the formula below:
(2) $\small S_{q,p;\Phi}=f_{q}(q,\Phi_{q})^{T}f_{p}(p;\Phi_{p}),\ p\in P,$
where $\Phi=[\Phi_{q},\Phi_{p}]$ denotes the retriever question and passage encoder parameters. We choose the final Top-k passage as below:
(3) $\small P_{R}=R(P,q)=\{p_{1},...,p_{k}\}=Top_{k}(min(S_{q,p;\Phi},p\in P)).$
The retrieved result in $P_{R}$ will be used as input and tuning material for knowledge integration, which can provide LLMs with more accurate knowledge.
@@ 4.2. Knowledge Integration
From the previous part, we acquired high-quality knowledge utilizing the retriever. To integrate this high-quality knowledge into the parameters of our model, tuning is an effective way. However, the original instruction tuning is inadequate for the OpenQA task, which is attributed to its limited scope of only aiding LLMs in comprehending the objectives of different tasks. Hence the knowledge integration $A_{K}=K(q,P)$ aims to enable the model to receive the necessary knowledge for generating the target answer. The detailed demonstration is clearly illustrated in Figure 2 part A.
Inspired by how pretraining enables LLMs to grasp knowledge (brown2020language, ; du2021glm, ; touvron2023llama, ), we also introduce the following tuning optimization to enable the model to learn more domain knowledge precisely. However, as one of the innovations of this paper, our Knowledge Integration module is neither learning all knowledge extensively like pre-training nor providing knowledge only in the prompt like contextual learning. The fine-tuning paradigm designed in this paper integrates retrieved refined knowledge into LLM through fine-tuning, ensuring the accuracy and domain specificity of the knowledge provided for LLM.
Specifically, tuning an LLM on domain knowledge $P$ is defined as follows:
(4) $\small\mathcal{L}_{r}=max_{\Phi}\sum_{p\in P}\sum_{t=1}^{|p|}log(P_{\Phi}(p_{t}|p_{<t})).$
@@
$max_{\Phi_{L}}(\lambda_{1}\sum_{p\in P}\sum_{t=1}^{|p|}log(P_{\Phi+\Phi_{L}}(p_{t}|p_{<t}))+\lambda_{2}\mathcal{L}_{f},$
Figure 3. The real cases of outputs after Knowledge Integration and Controllable Generation structure.
where $\mathcal{L}_{f}$ is the loss of original instruction tuning, $\Phi_{L}$ is the LoRA parameters and we only update LoRA parameters during the training process. The prompts we used are shown in Figure 2. In the experiment, we set $\lambda_{1}>\lambda_{2}>0$ to make the model focus on knowledge without losing the ability to answer questions.
However, due to the occasional presence of inaccuracies in the initial retrieval results $K_{R}$ leading mistake of $A_{K_{R}}=K(q,P_{R})$, we adopt this module with the input of either full passage $P$ or retrieved knowledge $P_{R}$ using $A_{K}=K(q,P)$ or $A_{K_{R}}=K(q,P_{R})$. These output options are presented to the next controllable generation step, allowing for a selective choice.
Incorporating knowledge during finetuning using this method leverages the advantages of the retriever, enabling the model to learn more precise and specialized knowledge. By storing domain knowledge in LLMs’ parameters, our work leads the model to learn knowledge more stably, resulting in superior performance. This method is used in conjunction with prompts using retrieved results to further enhance the model’s capability on answering questions.
@@ 4.3. Controllable Generation
From the knowledge integration model, we gain a draft of the answer. While the draft of this answer already contains all the knowledge needed to address the question, it still does not meet our goal of formatted generation. In order to align the model’s output with the dataset’s answers and our expectations, we need to introduce another framework for post-processing the answers from the last module, formally defined as $C(A_{K},F)=A$.
@@
The approach employed for achieving controllable generation utilizes the identical loss function as instruction tuning. However, due to the fact that the preceding step has yielded answers imbued with high-quality knowledge, aligning them into the format we require becomes a more manageable task than instruction tuning for the model to undertake. After fine-tuning the model with a certain amount of training data, we obtain the post-processing model $PostP$. The instruction we used for fine-tuning is demonstrated in part B in Figure 2.
@@
Additionally, derived from the previous step, we have acquired two distinct answers denoted as $PostP(A_{K},F)$ and $PostP(A_{K_{R}},F)$. The imminent task involves the meticulous selection of one among these to serve as the ultimate output. In order to make our answer reliable and controllable, we combine a reward model to effectively select answers that match the output style of the dataset and an external judging model when these two answers are severely different. The pair-wise loss of the reward model chooses the best answer between two answers, formulated as below:
(6) $\small\mathcal{L}_{R_{m}}=-log(\sigma(R_{m}(A_{+},F)-R_{m}(A_{-},F)),$
@@
where $R_{m}(A)$ is the average score of the embedding of the final token of A. However, relying solely on the reward model ensures that the model aligns with the preference for the target format F. To ensure coherence and consistency in the format of the output A, we need to introduce an additional model for analysis and filtering.
Specifically, we then design the judgment criteria score $S$ as below, encouraging answer format alignment and coherence enhancement when answers are long and similar:
@@
where $Q$ denotes questions, $A_{1}$ and $A_{2}$ respectively represents $PostP(A_{K})$ and $PostP(A_{K_{R}})$. $len(A)$ is the length of answers and $\overline{len(A)}$ is the mean of $A$, In this example, $A=\{A_{1},A_{2}\}$. Operations like $Exp$ aim to make $C_{S}$ score and $R_{m}$ score at the same scale, while $len(A)$ is designed to make the model focus more on fluency when the answer is longer. We adopt normalized inverse sentence frequency (NISF)(ke-etal-2022-ctrleval, ) to calculate answer consistency $C_{S}$, which is formally calculated as below:
(8) $\small C_{s}(QA)=NISF(A)logP_{\theta}(A|Q[M])\\ +NISF(Q)logP_{\theta}(Q|[M]A),$
(9) $\small NISF(Y_{j})=\frac{ISF(Y_{j})}{\Sigma_{k=1}^{M}ISF(Y_{k})},$
where $Q$, $\{Y_{1},...,Y_{j}\}$ denotes the question and the sentences of the question-answer pair. $[M]$ demotes masking when calculating probability using LM and $A$ denotes the answer. $ISF(Y)$ the inverse sentence frequency, which is related to the maximum of maximum inverse word frequency $(IWF)$ below:
(10) $\small IWF(\omega)=\frac{log(1+|C|)}{f_{\omega}},$
(11) $\small ISF(Y_{j})=max_{\omega\in Y_{j}}IWF(\omega),$
where $\{\omega_{11},...,\omega_{1i},...,\omega_{ji}\}$ denotes the words of the sentences. $|C|$ indicates the total number of sentences, $f_{\omega}$ denotes the frequency of $\omega$. After computing the score, we can employ the reward model to ensure answer coherence and format alignment or prioritize the correctness and reliability of answers derived through $Ext$, an external selection using chatGPT. Despite chatGPT’s performance not exceeding our results from the previous stage, it plays a role in augmenting overall capability. Details are presented in Section 6.3. The overall algorithm is delineated in Algorithm  1.
Algorithm 1 Overall pseudo code of GenKI
1:Knowledge base $K$, Retriever $R$, LLM $L$, question set $Q$
2:Answer $A$
3:Train $L_{1}$ and $L_{2}$ with $K$ and $K_{R}=R(K,Q)$ respectively
4$\triangleright$ Prompt I, II is used here
5:Gain $P_{3}$ using $L_{2}$ answering training queries
6:Train $L_{3}$ with $P_{3}$              $\triangleright$ Prompt III is used here
7:$A_{K}=L_{1}(Q),A_{K_{R}}=L_{2}(Q)$
8:$A_{1}=PostP(A_{K},F)=L_{3}(A_{K},F),A_{2}=PostP(A_{K_{R}},F)=L_{3}(A_{K_{R}},F)$
9:$S=Exp(|C_{s}(A_{1})-C_{s}(A_{2})|)-(|R_{m}(A_{1})-R_{m}(A_{2})|)/\overline{R_{m}(A)}*\overline{len(A)}$
10:if S<0 then
11:     $A=max_{A}(Rm(PostP(A_{K},{F}),F),$
12:                  $Rm(PostP(A_{K_{R}},{F}),F))$
13:else
14:     $A=Ext(PostP(A_{K},{F}),PostP(A_{K_{R}},F))$
15:     $\triangleright$ Prompt IV is used here
16:end if
By combining instruction-tuned LLM and a controllable answer selection model, we will gain an answer with both fluency brought by the reward model and answer-format accuracy brought by instruction-tuned LLM. Finally, to provide a clearer demonstration of how different components of GenKI contribute to its performance, we present real cases of GenKI’s outputs at different stages in Figure 3. This visualization allows us to observe GenKI’s transformation and refinement of outputs as it progresses through each stage of the process.
@@ 5.1. Datasets
Specifically, TriviaQA is a reading comprehension dataset containing over 650K training question-answer-evidence triples. The answers are mostly composed of entities in Wikipedia, containing one or two words. MSMARCO dataset contains 8,841,823 passages from documents retrieved by the Bing search engine, which provides the necessary information for curating the natural language answers. Different from TriviaQA, these answers belong to long and fluent sentences. CMRC-2018 is a dataset for Chinese Machine Reading Comprehension. It is a span-extraction reading comprehension dataset, which demands answers completely extracted from the original text. Following (dua-etal-2023-adapt, ), we processed reading comprehension datasets (CMRC and MSMARCO) by shuffling their reference, aligning the formal definition of the OpenQA, where knowledge must be searched from the knowledge base, not given.
@@ 5.2. Baselines
Dataset Format Length QA pairs
@@ 6. Results and Analysis
Baseline Type Method EMF1%
Task Specific Baselines GLaM-62B 75.8 -
@@
    $RA$: retriever-augmented baselines, $GA$: generator-augmented baselines (finetuned-LLM baselines), $B$: base model, $G$: GenKI
@@ 6.1. Overall Performance
    $RA$: retriever-augmented baselines, $GA$: generator-augmented baselines (finetuned-LLM baselines), $B$: base model, $G$: GenKI
@@ 6.1.1. Can GenKI work when the entire knowledge base is inaccessible?
In practical academic and industry contexts, models frequently depend on search engines to gather requisite knowledge, facing limitations in leveraging the entire knowledge base for fine-tuning. In such instances, our model ($X+K_{R}+PostP$) consistently surpasses baseline performances, demonstrating superiority by 3.3% in TriviaQA, 0.8% in MSMARCO, and 4.9% in CMRC. The comparative results are illustrated in Figure 5.

### v2

@@ 4. Methodology
To address the knowledge deficiency and uncontrollable generation issues in LLMs for OpenQA, our system combines three modules as illustrated in Fig. 2. On the whole, we decompose the model $A=M(q_{i},F,P)$ into two stage to ensure that the model focuses on either knowledge integration or controlled generation in each instance, avoiding the distributional differences brought by these two tasks. Specifically, we first train a knowledge retriever module $R(P,q_{i})=P_{R}$ to retrieve knowledge according to the query supplying the model during knowledge integration steps. Subsequently, once the retrieved knowledge is synthesized, a knowledge integration module $K(q_{i},P_{R})=A_{K}$ (Part A in Fig. 2) is deployed to generate answers by integrating the retrieved knowledge. Finally, a controllable generation module involves using a fine-tuned LLM on the target dataset format for post-processing the results, which is denoted as $\mathrm{PostP}$ and a Reward model $\mathrm{Rm}$ to further ensure that the output of our method is controlled. The function of each module will be introduced in detail below.
Figure 2. Overall framework of GenKI. The left part presents several examples of our prompts used in the knowledge integration and controllable generation module. The right part outlines the process of tuning LLMs to learn knowledge and answer questions in specific task situations.
@@ 4.1. Knowledge retriever
Firstly, given vast knowledge repositories, we should prioritize enabling our model to learn from them. However, feeding all the sentences $P$ from the knowledge base into the LLM may potentially result in the model’s inability to grasp essential focal points. Therefore, we need a retriever $R(P,q)=P_{R}\subset P$ to retrieve knowledge $P_{R}$ as the initial input provided for the next knowledge integration module.
Specifically, we use a Dense Passage Retriever (DPR) to retrieve knowledge from the sentence knowledge base $P$ according to the question. Setting the encoder as $f$, and the trainable parameters as $\Phi$, DPR computes question-sentence similarity $S$ for a given question $q$ and a set of sentences $P$ using the formula below:
(2) $S_{q,\ p,\ \Phi}=f_{q}(q,\Phi_{q})^{\mathrm{T}}f_{p}(p;\Phi_{p}),\ p\in P$
where $\Phi=[\Phi_{q},\Phi_{p}]$ denotes the retriever question and passage encoder parameters. We choose the final Top-k passage $P_{R}$ as below:
(3) $P_{R}=R(P,q)=\mathrm{Top}_{k}(\mathrm{min}\ (S_{q,\ p,\ \Phi},p\in P))$
The retrieved result in $P_{R}$ will be used as the input and tuning material for knowledge integration, which can provide LLMs with more accurate knowledge.
@@ 4.2. Knowledge integration
From the previous part, we acquire high-quality knowledge utilizing the retriever. To integrate this high-quality knowledge into the parameters of our model, tuning is an effective way. However, the original instruction tuning is inadequate for the OpenQA task, which is attributed to its limited scope of only aiding LLMs in comprehending the objectives of different tasks. Hence the knowledge integration $A_{K}=K(q_{i},P)$ aims to enable the model to receive the necessary knowledge for generating the target answer. The detailed demonstration is clearly illustrated in Fig. 2. Inspired by how pretraining enables LLMs to grasp knowledge (brown2020language, ; du2021glm, ; touvron2023llama, ), we introduce the following tuning optimization to enable the model to learn more domain knowledge precisely. However, as one of the innovations of this paper, our knowledge integration module is neither learning all knowledge extensively like pre-training nor providing knowledge only in the prompt like contextual learning. The fine-tuning paradigm designed in this paper integrates retrieved refined knowledge into LLM through fine-tuning, ensuring the accuracy and domain specificity of the knowledge provided for LLM.
Specifically, the tuning loss $\mathcal{L}_{r}$ of an LLM on domain passage $P$ is defined as follows:
(4) $\mathcal{L}_{r}=\mathop{\mathrm{max}}_{\Phi}\sum_{p_{i}\in P}\sum_{t=1}^{|p_{i}|}\mathrm{log\ }(\mathop{\mathrm{Prob\ }}_{\Phi}(p_{i}[t]|p_{i}[<t]))$
where || denotes the nomber of tokens of $p_{i}$.
@@
$\mathop{\mathrm{max\ }}_{\Phi_{L}}(\lambda_{1}\sum_{p\in P}\sum_{t=1}^{|p_{i}|}\mathrm{log\ }(\mathop{\mathrm{{Prob\ }}}_{\Phi+\Phi_{L}}(p_{i}[t]|p_{i}[<t])))+\lambda_{2}\mathcal{L}_{f},$
Figure 3. Real cases of outputs after Knowledge integration and controllable generation structure.
where $\mathcal{L}_{f}$ is the loss of original instruction tuning, $\Phi_{L}$ is the LoRA parameters and we only update LoRA parameters during the training process. The prompts we used are shown in Fig. 2. In the experiment, we set $\lambda_{1}>\lambda_{2}>0$ to make the model focus on knowledge without losing the ability to answer questions.
Given that retrieval can occasionally result in inaccuracies, leading to errors in $A^{\prime}=K(q,P_{R})$, where $A^{\prime}$ is the answer adopting retrieved passages $P_{R}=R(P,q_{i})$, it is necessary to retain the option of using the full passage $P$ without retrieval. Therefore, we utilize this module with inputs either from the complete passage $P$ or from the retrieved knowledge $P_{R}$, using $A=K(q,P)$ or $A^{\prime}=K(q,P_{R})$, respectively. These output options are then presented to the subsequent controllable generation step, allowing for a selective choice.
Incorporating knowledge during finetuning using this method leverages the advantages of the retriever, enabling the model to learn more precise and specialized knowledge. By storing domain knowledge in LLMs’ parameters, our work leads the model to learn knowledge more stably, resulting in superior performance.
@@ 4.3. Controllable generation
From the knowledge integration model, we gain a draft of the answer. While the draft of this answer already contains all the knowledge needed to address the question, it still does not meet our goal of formatted generation. In order to align the model’s output with the dataset’s answers and our expectations, we need to introduce another framework for post-processing the answers from the last module.
@@
The approach employed for achieving controllable generation utilizes the identical loss function as instruction tuning. However, due to the fact that the preceding step has yielded answers imbued with high-quality knowledge, aligning them into the format we require becomes a more manageable task than instruction tuning for the model to undertake. After fine-tuning the model with a certain amount of training data, we obtain the post-processing model $\mathrm{PostP}$. The instruction we used for fine-tuning is demonstrated in Part B in Fig. 2.
@@
Additionally, derived from the previous step, we have acquired two distinct answers, denoted as $\mathrm{PostP}(A_{K},F)$ and $\mathrm{PostP}(A_{K_{R}},F)$. The imminent task involves the meticulous selection of one among these to serve as the ultimate output. In order to make our answer reliable and controllable, we combine a reward model to effectively select answers that match the output style of the dataset and an external judging model when these two answers are severely different. The pair-wise loss $\mathcal{L}_{R_{m}}$ of the reward model chooses the best answer between two answers, formulated as below:
(6) $\mathcal{L}_{Rm}=-\mathrm{log\ }(\sigma(Rm(A_{+},F)-Rm(A_{-},F))$
@@
where $Rm(A)$ is the average score of the embedding of the final token of $A$, $A_{+}$ and $A_{-}$ denotes the positive and negative answer respectively. However, relying solely on the reward model ensures that the model aligns with the preference for the target format $F$. To ensure coherence and consistency in the format of the output $A$, we need to introduce an additional model for analysis and filtering. Specifically, we then design the judgment criteria score $S_{c}$ as below, encouraging answer format alignment and coherence enhancement when answers are long and similar:
(7) $\begin{aligned} S_{c}=&\mathrm{exp\ }(|C_{s}(QA_{1})-C_{s}(QA_{2})|)-\\ &\frac{|R_{m}(A_{1},F)-R_{m}(A_{2},F)|}{\overline{R_{m}(A,F)}\overline{\mathrm{len\ }(A)}})\end{aligned},$
where $Q$ denotes questions, $A_{1}$ and $A_{2}$ represents $\mathrm{PostP}(A_{K})$ and $\mathrm{PostP}(A_{K_{R}})$ respectively. $\mathrm{len\ }(A)$ is the length of answers and $\overline{\mathrm{len\ }(A)}$ is the mean of $A$, In this example, $A=\{A_{1},A_{2}\}$. Operations like $\mathrm{exp}(\cdot)=e^{\cdot}$ aim to make consistency score $C_{S}$ and reward score $Rm$ at the same scale, while $\mathrm{len\ }(A)$ is designed to make the model focus more on fluency when the answer is longer. We adopt Normalized Inverse Sentence Frequency (NISF)(ke-etal-2022-ctrleval, ) to calculate answer consistency $C_{S}$, which is formally calculated as below:
(8) $C_{s}(Q,A)=$ $\mathrm{NISF\ }(A)\mathrm{log\ }(\mathrm{Prob}_{\theta}(A|Q[\mathrm{Mask}]))$
$+\mathrm{NISF\ }(Q)\mathrm{log\ }(\mathrm{Prob}_{\theta}(Q|[\mathrm{Mask}]A))$
(9) $\mathrm{NISF\ }(Y_{j})=\frac{\mathrm{ISF}(Y_{j})}{\Sigma_{k=1}^{M}\mathrm{ISF}(Y_{k})},$
where $Q$ and $Y=\{Y_{1},Y_{2},\ldots,Y_{j}\}$ denote the question and the sentences of the question-answer pair, $\mathrm{Mask}]$ denotes masking to calculate probability using LM and $A$ denotes the answer. $\mathrm{ISF\ }(Y)$ is the inverse sentence frequency of $Y$, which is related to the maximum of maximum inverse word frequency $(\mathrm{IWF})$ below,
(10) $\mathrm{IWF}(\omega)=\frac{\mathrm{log\ }(1+\mathrm{Count})}{f_{\omega}}$
(11) $\mathrm{ISF}(Y_{j})=\mathop{\mathrm{max}}_{\omega\in Y_{j}}\mathrm{IWF}(\omega)$
where $\{\omega_{11},...,\omega_{1i},...,\omega_{ji}\}$ denotes the words of the sentences. $\mathrm{Count}$ indicates the total number of sentences, $f_{\omega}$ denotes the frequency of $\omega$. After computing the score, we can employ the reward model to ensure answer coherence and format alignment or prioritize the correctness and reliability of answers derived through $\mathrm{Ext}$, an external selection using chatGPT. Despite chatGPT’s performance not exceeding our results from the previous stage, it plays a role in augmenting overall capability. Details are presented in Section 6.3. The overall algorithm is delineated in Algorithm  1.
Algorithm 1 Overall pseudo code of GenKI
1:Knowledge base $K$, Retriever $R$, LLM $L_{1},\ L_{2},\ L_{3}$, question set $Q$
2:Answer AA
3:Train LLM $L_{1}$ and $L_{2}$ with $K$ and $K_{R}=R(K,Q)$ respectively $\triangleright$ Prompt I, II is used here
4:Gain format passage $P_{f}$ using $L_{2}$ answering training queries
5:Train LLM $L_{3}$ with $P_{f}$ $\triangleright$ Prompt III is used here
@@
[…]
9:if $S<0$ then
10:  $A=\mathop{\mathrm{max}}_{A}(\mathrm{Rm}(\mathrm{PostP}(A_{K},{F})),\mathrm{Rm}(\mathrm{PostP}(A_{K_{R}},{F})));$
11:else
12:  $A=\mathrm{Ext\ }(\mathrm{PostP}(A_{K},{F}),\mathrm{PostP}(A_{K_{R}},F));$
$\triangleright$ Prompt IV is used here
13:end if
By combining instruction-tuned LLM and a controllable answer selection model, we gain an answer with both fluency brought by the reward model and answer-format accuracy brought by instruction-tuned LLM. Finally, to provide a clearer demonstration of how different components of GenKI contribute to its performance, we present real cases of GenKI’s outputs at different stages in Fig. 3. This visualization allows us to observe GenKI’s transformation and refinement of outputs as it progresses through each stage of the process.
@@ 5.1. Datasets
Specifically, TriviaQA is a reading comprehension dataset containing over 650, 000 training question-answer-evidence triples. The answers are mostly composed of entities in Wikipedia, containing one or two words. MSMARCO dataset contains 8, 841, 823 passages from documents retrieved by the Bing search engine, which provides the necessary information for curating the natural language answers. Different from TriviaQA, these answers belong to long and fluent sentences. CMRC-2018 is a dataset for Chinese Machine Reading Comprehension. It is a span-extraction reading comprehension dataset, which demands answers completely extracted from the original text. Following (dua-etal-2023-adapt, ), we processed reading comprehension datasets (CMRC and MSMARCO) by shuffling their reference, aligning the formal definition of the OpenQA, where knowledge must be searched from the knowledge base, not given.
@@ 5.2. Baselines
Dataset Format Answer type QA pairs
@@ 6. Results and Analysis
Baseline type Method EM% F1%
Task
specific
baselines
GLaM-62B 75.8 -
@@
    RARA: retriever-augmented baselines, $GA$: generator-augmented baselines (finetuned-LLM baselines), $B$: base model, $G$: GenKI
@@ 6.1.1. Can GenKI work when the entire knowledge base is inaccessible?
In practical academic and industry contexts, models frequently depend on search engines to gather requisite knowledge, facing limitations in leveraging the entire knowledge base for fine-tuning. In such instances, our model ($X+K_{R}+\mathrm{PostP}$) consistently surpasses baseline performances, demonstrating superiority by 3.3% in TriviaQA, 0.8% in MSMARCO, and 4.9% in CMRC. The comparative results are illustrated in Figure 5.